# The Relative Entropy of Expectation and Price

### v6

@@ Abstract
for the price $p$ and payoff $P$ of a funded and hedged derivative security, where the price measure $\mathbb{E}$ has minimum entropy relative to economic expectations, and the parameter $\alpha$ matches the risk aversion of the investor. This price principle is easily applied to standard models for market evolution, with applications considered here in model risk analysis, deep hedging and quantum information.
@@ I Introduction
$-c_{2}\leq p\leq c_{1}$ (2)
The capital $c_{1}$ deployed by the first counterparty is partitioned into the variation margin $p$ and the initial margin $m_{1}:=c_{1}-p$, and the capital $c_{2}$ deployed by the second counterparty is partitioned into the variation margin $-p$ and the initial margin $m_{2}:=c_{2}+p$. With the initial margin deposited in the margin account, the final margin available to settle the derivative is $m_{1}(1+db/b)$ for the first counterparty and $m_{2}(1+db/b)$ for the second counterparty. This is sufficient to settle the contractual commitments of the derivative when:
$-m_{1}(1+db/b)\leq p+dp\leq m_{2}(1+db/b)$ (3)
@@
demonstrating that the discounted return on capital matches the discounted return on the derivative adjusted by a put option on $d(p/b)$ with strike $-c_{1}/b$, exercised when the first counterparty defaults, and a call option on $d(p/b)$ with strike $c_{2}/b$, exercised when the second counterparty defaults. When assessing the viability of the trade, it is these net returns that must be considered, accounting for funding costs and the possibility that the settling counterparty defaults on its commitments.
@@
In this essay, entropic foundations for pricing are investigated that do not rely on unrealistic assumptions regarding market completeness or the effectiveness of the trading strategy, properly account for residual model risk, and establish consistency across diverse applications including algorithmic trading, derivative pricing and hedging, and the modelling of margin and capital. In developing these foundations, the entropic risk metric has dual application: as the a priori measure of performance used to optimise the investment strategy; and as the determinant of margin used to protect against default. These results leverage the remarkable versatility and tractability of relative entropy as a universal metric of uncertainty, applicable across a wide range of models and algorithms.
@@ II.1 Entropic risk metric
[…]
@@
$\frac{d\mathbb{E}^{\alpha}}{d\mathbb{E}\hfill}:=\frac{\exp[-\alpha\,dc]}{\mathbb{E}[\exp[-\alpha\,dc]]}$ (17)
@@ II.2 Optimising entropic risk
Introduce the families of equivalent measures $\mathbb{E}^{\phi}$ and $\mathbb{E}^{\phi,\zeta}$ with Radon-Nikodym weights:
$\frac{d\mathbb{E}^{\phi}}{d\mathbb{E}\hfill}$ $:=\frac{\exp[-\phi\cdot dq]}{\mathbb{E}[\exp[-\phi\cdot dq]}$ (24)
$\frac{d\mathbb{E}^{\phi,\zeta}}{d\mathbb{E}\hfill}$ $:=\frac{\exp[-\phi\cdot dq-\frac{1}{2}dq\cdot\zeta\,dq]}{\mathbb{E}[\exp[-\phi\cdot dq-\frac{1}{2}dq\cdot\zeta\,dq]]}$
parametrised by the vector $\phi$ and the positive-definite matrix $\zeta$. Price measures that are equivalent to the expectation measure are identified within this collection from the equilibrium of the entropic-risk optimisation strategy. The parameter $\phi$ arises from the change of measure in the definition of the entropic risk metric, which depends on the existence of the cumulant generating function. If this does not exist, the quadratic-regularised measure with parameter $\zeta$ is used instead. Regularisation is removed in the limit $\zeta\to 0$; if the strategy does not survive this limit, then regularisation, which moderates the impact of large events, is retained as a parameter of the model. In the following, the cumulant generating function is assumed to exist, implying the use of quadratic regularisation wherever necessary.
@@
$0=\mathbb{E}^{\phi_{1}\oplus\cdots\oplus\phi_{n}}[dq_{i}]$ (30)
for $i=1,\ldots,n$. This provides $n$ conditions for the $n$ components of the optimal portfolio. While these calibration conditions need to be solved simultaneously, market tiering is exploited to incrementally build the optimal portfolio from the independent contributions of sector investors, localising at each tier the risk management activity for the corresponding market expert. In this perspective, there are $n(n+1)/2$ calibration conditions for $n$ sub-markets:
00 0 $=\mathbb{E}^{\phi_{1}^{1}}[dq_{1}]$ (31)
00 0 $=\mathbb{E}^{\phi_{1}^{2}\oplus\phi_{2}^{2}}[dq_{1}]=\mathbb{E}^{\phi_{1}^{2}\oplus\phi_{2}^{2}}[dq_{2}]$
@@
00 0 $=\mathbb{E}^{\phi_{1}^{n}\oplus\cdots\oplus\phi_{n}^{n}}[dq_{1}]=\cdots=\mathbb{E}^{\phi_{1}^{n}\oplus\cdots\oplus\phi_{n}^{n}}[dq_{n}]$
where the calibration conditions in the $i$th row maximise the entropy-adjusted mean of the $i$th optimal portfolio $\phi^{i}=\phi_{1}^{i}\oplus\cdots\oplus\phi_{i}^{i}$ constituted in the lowest $i$ market tiers only. The $i$th investor then takes responsibility for determining the primary component $\psi_{i}$ of the $i$th optimal portfolio and its delta $\delta^{i}=\delta_{1}^{i}\oplus\cdots\oplus\delta_{i-1}^{i}$ to the $(i-1)$th optimal portfolio in the underlying sub-market, using the invertible linear change of variables:
@@
$\frac{d\mathbb{E}^{i}}{d\mathbb{E}\hfill}:=\frac{\exp[-(\phi_{1}^{i}\cdot dq_{1}+\cdots+\phi_{i}^{i}\cdot dq_{i})]}{\mathbb{E}[\exp[-(\phi_{1}^{i}\cdot dq_{1}+\cdots+\phi_{i}^{i}\cdot dq_{i})]]}$ (33)
@@ III Entropic pricing and hedging
Underlying only: In this strategy, the investment comprises an underlying portfolio $\omega$ satisfying the initial condition: $0=\omega\cdot q$ (36) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[\omega\cdot dq]=$ (37) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha\omega\cdot dq]$ at the risk aversion $\alpha>0$. The optimal portfolio $\omega$ that maximises this satisfies the calibration condition: $0=\mathbb{E}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (38) where $r\,dt$ is the Lagrange multiplier used to implement the self-funding condition. Underlying and derivative: In this strategy, the investment comprises the derivative security and an underlying portfolio $(\omega-\delta)$ satisfying the initial condition: $0=p+(\omega-\delta)\cdot q$ (39) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[dp+(\omega-\delta)\cdot dq]=$ (40) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]$ at the risk aversion $\alpha>0$. The hedge portfolio $\delta$ that maximises this satisfies the hedge condition: $0=\mathbb{E}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]]$ (41) where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition.
@@
These two strategies offer different entropy-adjusted mean returns to the investor. When the derivative security has no intrinsic value beyond its contractual link with the underlying securities, the efficient market equilibrates to the indifferent state wherein these entropy-adjusted means align, leading to the price condition: $\mathbb{E}[\alpha][\omega\cdot dq]=\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$ (42) The calibration condition is solved to derive the optimal portfolio $\omega$ of underlying securities. The hedge and price conditions are then solved simultaneously to derive the initial price $p$ from the final price $P$ of the derivative security and the hedge portfolio $\delta$ of underlying securities that re-establishes optimality for the combined portfolio.
@@
The market of underlying securities accessible to the investor performs the two key functions of funding and hedging for the derivative security. Fix a finite investment horizon with initial time $t$ and final time $T=t+dt$, and assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $Q=q+dq$ and a derivative security with initial price $p$ and final price $P=p+dp$. The investor quantifies a priori the performance of the investment strategy using an initial assessment of the final economic state, encapsulated in the expectation measure $\mathbb{E}$ that maps final variables to their initial expectations. Following the strategy developed in the previous section, the optimal portfolio then maximises the entropy-adjusted mean return at a specified risk aversion $\alpha$ for the investor.
As summarised in figure 5, there are two self-funded investment strategies to consider: one that includes only underlying securities; and one that also includes the derivative security. In the first strategy, the portfolio $\omega$ of underlying securities is optimal when it maximises the entropy-adjusted mean return $\mathbb{E}[\alpha][\omega\cdot dq]$. The optimal portfolio thus satisfies the calibration condition:
$0=\mathbb{E}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (43)
@@
where $r\,dt$ is the Lagrange multiplier used to implement the self-funding condition $\omega\cdot q=0$. Adjusting for the opportunity and risk that adding the derivative security presents, in the second strategy the optimal portfolio $\omega$ is modified by the hedge portfolio $\delta$, and optimality is restored in the combined portfolio when it maximises the entropy-adjusted mean return $\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$. The hedge portfolio thus satisfies the hedge condition:
$0=\mathbb{E}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]]$ (44)
@@
the derivative security neither advantages nor disadvantages the investment strategy.
The calibration, hedge and price conditions, together with the two self-funding conditions, are the foundations of entropic pricing. By avoiding the unrealistic assumptions that enable risk-neutral theory, entropic pricing can be consistently applied to pricing and hedging across a range spanning everything from algorithmic trading to XVA. These applications are explored in more depth in the following sections.

### v7

@@ Abstract
for the price $p$ and payoff $P$ of a funded and hedged derivative security, where the price measure $\mathbb{E}$ has minimum entropy relative to economic expectations, and the parameter $\alpha$ matches the risk aversion of the investor. This price principle is easily applied to standard models for market evolution, with applications considered here in model risk analysis, data-driven hedging and quantum information.
@@ I Introduction
$c_{2}\leq p\leq c_{1}$ (2)
The capital $c_{1}$ deployed by the first counterparty is partitioned into the variation margin $p$ and the initial margin $m_{1}:=c_{1}-p$, and the capital $c_{2}$ deployed by the second counterparty is partitioned into the variation margin p-p and the initial margin $m_{2}:=c_{2}+p$. With the initial margin deposited in the margin account, the final margin available to settle the derivative is $m_{1}(1+db/b)$ for the first counterparty and $m_{2}(1+db/b)$ for the second counterparty. This is sufficient to settle the contractual commitments of the derivative when:
$m_{1}(1+db/b)\leq p+dp\leq m_{2}(1+db/b)$ (3)
@@
demonstrating that the discounted return on capital matches the discounted return on the derivative adjusted by a put option on $d(p/b)$ with strike $c_{1}/b$, exercised when the first counterparty defaults, and a call option on $d(p/b)$ with strike $c_{2}/b$, exercised when the second counterparty defaults. When assessing the viability of the trade, it is these net returns that must be considered, accounting for funding costs and the possibility that the settling counterparty defaults on its commitments.
@@
In this essay, entropic foundations for pricing are investigated that do not rely on unrealistic assumptions regarding market completeness or the effectiveness of the trading strategy, properly account for residual model risk, and establish consistency across diverse applications including algorithmic trading, derivative pricing and hedging, and the modelling of margin and capital. In developing these foundations, the entropic risk metric has dual application: as the measure of performance used to optimise the investment strategy; and as the determinant of margin used to protect against default. These results leverage the remarkable versatility and tractability of relative entropy as a universal metric of uncertainty, applicable across a wide range of models and algorithms.
@@ II.1 Entropic risk metric
$={}$ $\mathbb{E}[dc]-\frac{1}{2}\alpha\mathbb{V}[dc]+\sum_{n=3}^{\infty}\frac{(-\alpha)^{n-1}}{n!}\mathbb{M}_{n}[dc]$
@@
$\frac{\displaystyle d\mathbb{E}^{\alpha}}{\displaystyle d\mathbb{E}\hfill}:=\frac{\exp[-\alpha\,dc]}{\mathbb{E}[\exp[-\alpha\,dc]]}$ (17)
@@ II.2 Optimising entropic risk
Introduce the equivalent measure $\mathbb{E}^{\phi}$ parametrised by the vector $\phi$ with Radon-Nikodym weight:
$\frac{\displaystyle d\mathbb{E}^{\phi}}{\displaystyle d\mathbb{E}\hfill}:=\frac{\exp[-\phi\cdot dq]}{\mathbb{E}[\exp[-\phi\cdot dq]}$ (24)
Price measures that are equivalent to the expectation measure are identified within this family from the equilibrium of the entropic-risk optimisation strategy. The parameter $\phi$ arises from the change of measure in the definition of the entropic risk metric, which depends on the existence of the cumulant generating function. If this does not exist, the quadratic-regularised measure is used instead, with Radon-Nikodym weight:
$\frac{\displaystyle d\mathbb{E}^{\phi}}{\displaystyle d\mathbb{E}\hfill}:=\frac{\exp[-\phi\cdot dq-\frac{1}{2}dq\cdot\zeta\,dq]}{\mathbb{E}[\exp[-\phi\cdot dq-\frac{1}{2}dq\cdot\zeta\,dq]]}$ (25)
regularised by the positive-definite matrix $\zeta$. Regularisation is removed in the limit $\zeta\to 0$; if the strategy does not survive this limit, then regularisation, which moderates the impact of large events, is retained as a parameter of the model. In the following, the cumulant generating function is assumed to exist, implying the use of quadratic regularisation wherever necessary.
@@
$0=\mathbb{E}^{\phi_{1}\oplus\cdots\oplus\phi_{n}}[dq_{1}]=\cdots=\mathbb{E}^{\phi_{1}\oplus\cdots\oplus\phi_{n}}[dq_{n}]$ (31)
This provides $n$ conditions for the $n$ components of the optimal portfolio. While these calibration conditions need to be solved simultaneously, market tiering is exploited to incrementally build the optimal portfolio from the independent contributions of sector investors, localising at each tier the risk management activity for the corresponding market expert. In this perspective, there are $n(n+1)/2$ calibration conditions for $n$ sub-markets:
$0$ $=\mathbb{E}^{\phi_{1}^{1}}[dq_{1}]$ (32)
$0$ $=\mathbb{E}^{\phi_{1}^{2}\oplus\phi_{2}^{2}}[dq_{1}]=\mathbb{E}^{\phi_{1}^{2}\oplus\phi_{2}^{2}}[dq_{2}]$
@@
$0$ $=\mathbb{E}^{\phi_{1}^{n}\oplus\cdots\oplus\phi_{n}^{n}}[dq_{1}]=\cdots=\mathbb{E}^{\phi_{1}^{n}\oplus\cdots\oplus\phi_{n}^{n}}[dq_{n}]$
where the calibration conditions in the $i$th row maximise the entropy-adjusted mean of the $i$th optimal portfolio $\phi^{i}=\phi_{1}^{i}\oplus\cdots\oplus\phi_{i}^{i}$ constituted in the lowest $i$ market tiers only. The $i$th investor then takes responsibility for determining the primary component $\psi_{i}$ of the $i$th optimal portfolio and its delta $\delta^{i}=\delta_{1}^{i}\oplus\cdots\oplus\delta_{i-1}^{i}$ to the (i1)(i-1)th optimal portfolio in the underlying sub-market, using the invertible linear change of variables:
@@
$\frac{\displaystyle d\mathbb{E}^{i}}{\displaystyle d\mathbb{E}\hfill}:=\frac{\exp[-(\phi_{1}^{i}\cdot dq_{1}+\cdots+\phi_{i}^{i}\cdot dq_{i})]}{\mathbb{E}[\exp[-(\phi_{1}^{i}\cdot dq_{1}+\cdots+\phi_{i}^{i}\cdot dq_{i})]]}$ (34)
@@ III Entropic pricing and hedging
Underlying only: In this strategy, the investment comprises an underlying portfolio $\omega$ satisfying the initial condition: $\omega\cdot q=0$ (37) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[\omega\cdot dq]=$ (38) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha\omega\cdot dq]$ at the risk aversion $\alpha>0$. The optimal portfolio $\omega$ that maximises this satisfies the calibration condition: $\mathbb{E}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]=0$ (39) where rdtr\,dt is the Lagrange multiplier used to implement the self-funding condition. Underlying and derivative: In this strategy, the investment comprises the derivative security and an underlying portfolio $(\omega-\delta)$ satisfying the initial condition: $p+(\omega-\delta)\cdot q=0$ (40) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[dp+(\omega-\delta)\cdot dq]=$ (41) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]$ at the risk aversion $\alpha>0$. The hedge portfolio $\delta$ that maximises this satisfies the hedge condition: $\mathbb{E}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]]=0$ (42) where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition.
@@
These two strategies offer different entropy-adjusted mean returns to the investor. When the derivative security has no intrinsic value beyond its contractual link with the underlying securities, the efficient market equilibrates to the indifferent state with matching entropy-adjusted means, leading to the price condition: $\mathbb{E}[\alpha][\omega\cdot dq]=\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$ (43) The calibration condition is solved to derive the optimal portfolio $\omega$ of underlying securities, and the hedge condition is then solved to derive the hedge portfolio $\delta$ of underlying securities that re-establishes optimality for the combined portfolio. With these portfolios, the fair initial price $p$ of the derivative security is derived from the price condition.
Define the price measure $\mathbb{E}^{\phi}$ equivalent to the expectation measure $\mathbb{E}$ with Radon-Nikodym weight: $\frac{\displaystyle d\mathbb{E}^{\phi}}{\displaystyle d\mathbb{E}^{\vphantom{\phi}}\hfill}:=\frac{\displaystyle\exp[-\phi\cdot dq]}{\displaystyle\mathbb{E}^{\vphantom{\phi}}[\exp[-\phi\cdot dq]]}$ (44) While there is no explicit expression for the unit optimal portfolio $\phi:=\alpha\omega$ that solves the calibration condition, it is efficiently discovered by the Newton-Raphson scheme initialised with $\phi=0$ and repeating the step: $r\,dt:={}$ (45) $\hskip-28.45274pt\frac{\displaystyle(\phi+\mathbb{V}^{\phi}[dq]^{-1}\mathbb{E}^{\phi}[dq])\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\cdot q}$ $\partial\phi={}$ $\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{E}^{\phi}[dq]-qr\,dt)$ until the required accuracy is achieved. Define the price measure $\mathbb{E}^{\phi\delta}:=\mathbb{E}^{(\phi-\alpha\delta)\oplus\alpha}$ equivalent to the price measure $\mathbb{E}^{\phi}$ with Radon-Nikodym weight: $\frac{\displaystyle d\mathbb{E}^{\phi\delta}}{\displaystyle d\mathbb{E}^{\phi}\hfill}:=\frac{\displaystyle\exp[-\alpha(P-\delta\cdot dq)]}{\displaystyle\mathbb{E}^{\phi}[\exp[-\alpha(P-\delta\cdot dq)]]}$ (46) While there is no explicit expression for the hedge portfolio $\delta$ that solves the hedge and price conditions, it is efficiently discovered by the Newton-Raphson scheme initialised with $\delta=0$ and repeating the step: $(r+\alpha s)\,dt:={}$ (47) $\hskip-28.45274pt\frac{\displaystyle\alpha\mathbb{E}^{\phi}[\alpha][P-\delta\cdot Q]+\mathbb{V}^{\phi\delta}[dq]^{-1}\mathbb{E}^{\phi\delta}[dq]\cdot\mathbb{E}^{\phi\delta}[Q]}{\displaystyle\mathbb{V}^{\phi\delta}[dq]^{-1}q\cdot\mathbb{E}^{\phi\delta}[Q]}$ $\partial\delta={}$ $-\mathbb{V}^{\phi\delta}[dq]^{-1}(\mathbb{E}^{\phi\delta}[dq]-q(r+\alpha s)\,dt)/\alpha$ until the required accuracy is achieved.
@@
Entropic Pricing and Hedging: Normally Distributed Returns
Fix a finite investment horizon and let $\mathbb{E}$ be the expectation measure mapping variables at the final time $T=t+dt$ to their expectations at the initial time $t$. Assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $Q=q+dq$ and a derivative security with initial price $p$ and final price $P=p+dp$. Also assume that the underlying and derivative returns have joint normal distribution in the expectation measure. Consider two self-funded strategies: one comprising underlying securities only; and one that also includes the derivative security.
@@
Underlying only: In this strategy, the investment comprises an underlying portfolio $\omega$ satisfying the initial condition: $\omega\cdot q=0$ (48) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[\omega\cdot dq]=$ (49) $\mathbb{E}[\omega\cdot dq]-\frac{1}{2}\alpha\mathbb{V}[\omega\cdot dq]$ at the risk aversion $\alpha>0$. The optimal portfolio $\omega$ that maximises this is given by the calibration solution: $r\,dt$ $=\frac{\displaystyle\mathbb{V}[dq]^{-1}q\cdot\mathbb{E}[dq]}{\displaystyle\mathbb{V}[dq]^{-1}q\cdot q}$ (50) $\alpha\omega$ $=\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)$ where rdtr\,dt is the Lagrange multiplier used to implement the self-funding condition. Underlying and derivative: In this strategy, the investment comprises the derivative security and an underlying portfolio $(\omega-\delta)$ satisfying the initial condition: $p+(\omega-\delta)\cdot q=0$ (51) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[dp+(\omega-\delta)\cdot dq]=$ (52) $\mathbb{E}[dp+(\omega-\delta)\cdot dq]-\frac{1}{2}\alpha\mathbb{V}[dp+(\omega-\delta)\cdot dq]$ at the risk aversion $\alpha>0$. The hedge portfolio $\delta$ that maximises this is given by the hedge solution: $s\,dt$ $=\frac{\displaystyle p-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,dp]\cdot q}{\displaystyle\mathbb{V}[dq]^{-1}q\cdot q}$ (53) $\delta$ $=\mathbb{V}[dq]^{-1}(\mathbb{V}[dq,dp]+qs\,dt)$ where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition.
These two strategies offer different entropy-adjusted mean returns to the investor. When the derivative security has no intrinsic value beyond its contractual link with the underlying securities, the efficient market equilibrates to the indifferent state with matching entropy-adjusted means, leading to the price solution $p=\delta\cdot q$ with hedge portfolio: $\delta=\mathbb{V}[dq]^{-1}\left(\mathbb{V}[dq,P]+\frac{1}{\alpha}\left(\sqrt{1+2\alpha\frac{\displaystyle\mathbb{E}[\alpha][P-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot Q]}{\displaystyle\mathbb{V}[dq]^{-1}\bar{q}\cdot\bar{q}}}-1\right)\bar{q}\right)$ (54) where $\bar{q}:=q(1+r\,dt)$ inflates the initial underlying price $q$ by the implied funding rate $r$. The hedge portfolio of underlying securities simultaneously funds and hedges the derivative security, and the price solution decomposes the initial derivative price as the weighted sum of its component initial underlying prices.
Figure 6: When the underlying and derivative returns have joint normal distribution in the expectation measure, the calibration, hedge and price conditions are solved explicitly in terms of their means and covariances.
Entropic Pricing and Hedging: Continuous Price Diffusion
Fix a finite investment horizon and let $\mathbb{E}$ be the expectation measure mapping variables at the final time $T=t+dt$ to their expectations at the initial time $t$. Assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $Q=q+dq$ and a derivative security with initial price $p$ and final price $P=p+dp$. Also assume that the prices diffuse continuously, so that their covariances are $O[dt]$ and their higher cumulants are $O[dt^{2}]$. Consider two self-funded strategies: one comprising underlying securities only; and one that also includes the derivative security.
Underlying only: In this strategy, the investment comprises an underlying portfolio $\omega$ satisfying the initial condition: $\omega\cdot q=0$ (55) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[\omega\cdot dq]=$ (56) $\mathbb{E}[\omega\cdot dq]-\frac{1}{2}\alpha\mathbb{V}[\omega\cdot dq]+O[dt^{2}]$ at the risk aversion $\alpha>0$. The optimal portfolio $\omega$ that maximises this is given by the calibration solution: $r\,dt$ $=\frac{\displaystyle\mathbb{V}[dq]^{-1}q\cdot\mathbb{E}[dq]}{\displaystyle\mathbb{V}[dq]^{-1}q\cdot q}+O[dt^{2}]$ (57) $\alpha\omega$ $=\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)+O[dt]$ where rdtr\,dt is the Lagrange multiplier used to implement the self-funding condition. Underlying and derivative: In this strategy, the investment comprises the derivative security and an underlying portfolio $(\omega-\delta)$ satisfying the initial condition: $p+(\omega-\delta)\cdot q=0$ (58) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}$ $[\alpha][dp+(\omega-\delta)\cdot dq]=$ (59) $\mathbb{E}[dp+(\omega-\delta)\cdot dq]-\frac{1}{2}\alpha\mathbb{V}[dp+(\omega-\delta)\cdot dq]+O[dt^{2}]$ at the risk aversion $\alpha>0$. The hedge portfolio $\delta$ that maximises this is given by the hedge solution: $s\,dt$ $=\frac{\displaystyle p-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,dp]\cdot q}{\displaystyle\mathbb{V}[dq]^{-1}q\cdot q}+O[dt^{2}]$ (60) $\delta$ $=\mathbb{V}[dq]^{-1}(\mathbb{V}[dq,dp]+qs\,dt)+O[dt]$ where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition.
These two strategies offer different entropy-adjusted mean returns to the investor. When the derivative security has no intrinsic value beyond its contractual link with the underlying securities, the efficient market equilibrates to the indifferent state with matching entropy-adjusted means, leading to the price solution: $p=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[P]-\frac{1}{2}\alpha\mathbb{V}[P-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot dq]+O[dt^{2}]$ (61) The first term expresses the discounted derivative price as a martingale in the price measure. The second term then adjusts the price to account for the residual risk after hedging.
Figure 7: When the underlying and derivative prices diffuse continuously in the expectation measure, the calibration, hedge and price conditions are solved explicitly to $O[dt]$ in terms of their means and covariances. These solutions can be used when the portfolio is rebalanced continuously.
The market of underlying securities accessible to the investor performs the two key functions of funding and hedging for the derivative security. Fix a finite investment horizon with initial time $t$ and final time $T=t+dt$, and assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $Q=q+dq$ and a derivative security with initial price $p$ and final price $P=p+dp$. The investor quantifies the performance of the investment strategy using an initial assessment of the final economic state, encapsulated in the expectation measure $\mathbb{E}$ that maps final variables to their initial expectations. Following the strategy developed in the previous section, the optimal portfolio then maximises the entropy-adjusted mean return at a specified risk aversion $\alpha$ for the investor.
There are two self-funded investment strategies to consider: one that includes only underlying securities; and one that also includes the derivative security. In the first strategy, the portfolio $\omega$ of underlying securities is optimal when it maximises the entropy-adjusted mean return $\mathbb{E}[\alpha][\omega\cdot dq]$. The optimal portfolio thus satisfies the calibration condition:
$\mathbb{E}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]=0$ (62)
where rdtr\,dt is the Lagrange multiplier used to implement the self-funding condition $\omega\cdot q=0$. Adjusting for the opportunity and risk that adding the derivative security presents, in the second strategy the optimal portfolio $\omega$ is modified by the hedge portfolio $\delta$, and optimality is restored in the combined portfolio when it maximises the entropy-adjusted mean return $\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$. The hedge portfolio thus satisfies the hedge condition:
$\mathbb{E}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]]=0$ (63)
@@
the derivative security neither advantages nor disadvantages the investment strategy. Since it has no value beyond its contractual links, this indifference relationship is assumed as the equilibrium condition for the derivative price. The market mechanism that achieves equilibrium is considered in more detail later in the essay.
The calibration, hedge and price conditions, together with the two self-funding conditions, are the foundations of entropic pricing. By avoiding the unrealistic assumptions that enable risk-neutral theory, entropic pricing can be consistently applied to pricing and hedging across a range spanning everything from algorithmic trading to XVA and capital modelling.
III.1 Calibration
Investments are self-funded within the underlying market, and the entropic model does not assume the existence of a risk-free funding rate for investments that are not self-funded. Calibration identifies the price measure $\mathbb{E}^{\phi}$ equivalent to the expectation measure $\mathbb{E}$ with Radon-Nikodym weight:
$\frac{\displaystyle d\mathbb{E}^{\phi}}{\displaystyle d\mathbb{E}\hfill}:=\frac{\displaystyle\exp[-\phi\cdot dq]}{\displaystyle\mathbb{E}[\exp[-\phi\cdot dq]]}$ (65)
where the unit optimal portfolio $\phi$ and implied funding rate $r$ are solved from the calibration and self-funding conditions:
$qr\,dt$ $=\mathbb{E}^{\phi}[dq]$ (66)
$\phi\cdot q$ $=0$
The implied funding rate defined here originates in the optimisation as the Lagrange multiplier for the self-funding condition.
When the portfolio is rebalanced on the discrete schedule $t_{0}<t_{1}<\cdots<t_{n}$, at each time $t_{i}$ the unit optimal portfolio $\phi_{i}$ is determined at the start of the investment interval $t_{i}<t_{i+1}$ and the calibration condition is compounded to generate the price measure $\mathbb{E}_{i}^{\phi}$ equivalent to the expectation measure $\mathbb{E}_{i}$:
$\frac{\displaystyle d\mathbb{E}_{i}^{\phi}}{\displaystyle d\mathbb{E}_{i}\hfill}:={}$ $\prod_{j=i}^{n-1}\frac{\displaystyle\exp[-\phi_{j}\cdot dq_{j}]}{\displaystyle\mathbb{E}_{j}[\exp[-\phi_{j}\cdot dq_{j}]]}$ (67)
$={}$ $\exp\!\left[\sum_{j=i}^{n-1}(-\phi_{j}\cdot dq_{j}-\log\mathbb{E}_{j}\exp[-\phi_{j}\cdot dq_{j}])\right]$
for variables on the discrete schedule. In the limit of continuous rebalancing, the sum in the exponent is replaced by the integral.